OpenStudy (anonymous):

Prove the trig identity sinx(tanx*cosx-cotx*cosx)=1-2cos^2x.

OpenStudy (anonymous):

sinx*cosx(tanx-cotx)=1-2cos^2x sinx*cosx(sinx/cosx-cosx/sinx)=1-2cos^2x sinx*cosx[sin^2x/(cosx*sinx)-cos^2x/(sinx*cosx)]=1-2cos^2x sinx*cosx[(sin^2x-cos^2x)/(sinx*cosx)]=1-2cos^2x sin^2x-cos^2x=1-2cos^2x

OpenStudy (mertsj):

Are you done?

OpenStudy (anonymous):

No, then I kinda got stuck there.

OpenStudy (anonymous):

Did I get something wrong in the process?

OpenStudy (mertsj):

replace sin^2x with 1-cos^2x on the left side and you'll be done.

OpenStudy (anonymous):

Oh... I can't believe I didn't realize that... looks like I need to go back to familiarizing my fundamental trig identities. Thanks!

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

*notice